Hello Tutors of Algebra.com,
I don't if is the right section but here goes
A lab technician has a 40% alcohol solution and a 60% alcohol solution. He needs to make 100 mL of solution that is 48% alcohol . How much of each available solution should he use ?
thanks in advance :D

Answer by stanbon(75887)

A lab technician has a 40% alcohol solution and a 60% alcohol solution. He needs to make 100 mL of solution that is 48% alcohol . How much of each available solution should he use ?
--------
Equation:
alcohol + alcohol = alcohol
---------
0.40x + 0.60(100-x) = 0.48*100
---
Multiply thru by 100 to get:
40x + 60*100 - 60x = 48*100
-20x = -12*100
x = 60 mL (amt. of 40% solution needed)
100-x = 40mL (amt. of 60% solution needed)
============
Cheers,
Stan H.